OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the points (-4,2) and (1,-1).

OpenStudy (anonymous):

1st Find the Slope....and then use Point Slope Form

OpenStudy (anonymous):

This is going to require a few steps. The first step is to find the slope given the two points. You will use the formula for slope: \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

Y=-3/4X-1 thats what I had got

OpenStudy (anonymous):

I got this for slope: \[-\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

so is that what I should put down?

OpenStudy (anonymous):

Well once you have the slope you need to use the point slope formula: \[y-y _{1}=m(x-x _{1})\]